Please help im stuck on this one problem!
Y^2-2y-63=0

Answer by stanbon(75887)

Y^2-2y-63=0
----------------
y^2-9y+7y-63 = 0
---
y(y-9)+7(y-9) = 0
---
(y-9)(y+7) = 0
y = 9 or y = -7
====================
Cheers,
Stan H.
=========